OpenStudy (anonymous):

Solve x2 + 10x + 7 = 0 by completing the square. Which equation is used in the process?

OpenStudy (anonymous):

a^2+2ab+c^2 .... so now here 2ab=2*5*1 and now link a which is 1 so be is 5 so add some to make 7 to 25 add 18 both sides then (x+5)^2 you will get (x+5)^2=18

OpenStudy (anonymous):

do u understand the meaning here

OpenStudy (anonymous):

The correct procedure for "completing the square" is as follows: fir shift the constant to the other side x^2 + 10x = -7 Now add the square of half of the coefficient of x to both sides here coefficient of x is 10, its half is 5, so add 5^2 to both sides x^2 + 10x + 5^2 = -7 + 5^2 Now LHS become (x+5)^2, so we can write (x+5)^2 = -7 + 25 taking root of both sides x+5 = +/-root(18) x = -5 +/- 3root2 x = -5 + 3root2 or x = -5 - 3root2 this is the full n proper solution....

OpenStudy (anonymous):

yea but theres no need for value of x ... good way to explain

OpenStudy (anonymous):

thanks vinit :) the question said "solve by completing the square" so i guess we have to find values of x.... What do you say ????

OpenStudy (anonymous):

completing the square not finding the value of x ??

OpenStudy (anonymous):

what do u do when u "solve" an equation???

OpenStudy (anonymous):

By the way vinit, r u frm Maharashtra?????

OpenStudy (anonymous):

solve for wat u hav to solve if they ask u to write a then write a not abcd ... u kno dnt mind kid its ok may be ur right k newazz explanation i like it